OpenStudy (anonymous):

Help with FUNCTION NOTATIONS AND GRAPHS. Can someone tell me a little bit about this algebra lesson? I've watched videos and read lessons but I still do not understand what it is talking about. Here's a question: If g(x) = x2 + 2, find g(3). I think you just plug in the 3 for x. But I don't undertand the graphing. Can someone link some videos or just type a paragraph on function notations so that I can understand better? Thank you!

OpenStudy (anonymous):

There are two parts to this, and I think this is where your confusion is coming from. The first part is the graph of the function itself. If you graph \[g(x)=y=x^2+2\]it'll look something like this 

OpenStudy (anonymous):

But finding g(3) is a little different. You're trying to evaluate the function at a point.  Like you said, you substitute 3 for x, giving you the following: \[g(x)=x^2+2\]\[g(3)=(3)^2+2\]\[g(3)=11\] Basically, the function tells you the height of the graph at any particular x-value. You ask the function, how high is the graph at 3? 11. How high is the graph at 4? 18. If your mother ever tracked your height on a wall according to your age, it's basically the same thing.